### v2

# Fundamental computational limits of weak learnability in high-dimensional multi-index models

Abstract
Multi-index models — functions which only depend on the covariates through a non-linear transformation of their projection on a subspace — are a useful benchmark for investigating feature learning with neural networks. This paper examines the theoretical boundaries of efficient learnability in this hypothesis class, focusing particularly on the minimum sample complexity required for weakly recovering their low-dimensional structure with first-order iterative algorithms, in the high-dimensional regime where the number of samples is $n=\alpha d$ is proportional to the covariate dimension $d$. Our findings unfold in three parts: (i) first, we identify under which conditions a trivial subspace can be learned with a single step of a first-order algorithm for any $\alpha\!>\!0$; (ii) second, in the case where the trivial subspace is empty, we provide necessary and sufficient conditions for the existence of an easy subspace consisting of directions that can be learned only above a certain sample complexity $\alpha\!>\!\alpha_{c}$. The critical threshold $\alpha_{c}$ marks the presence of a computational phase transition, in the sense that it is conjectured that no efficient iterative algorithm can succeed for $\alpha\!<\!\alpha_{c}$. In a limited but interesting set of really hard directions —akin to the parity problem— $\alpha_{c}$ is found to diverge. Finally, (iii) we demonstrate that interactions between different directions can result in an intricate hierarchical learning phenomenon, where some directions can be learned sequentially when coupled to easier ones. Our analytical approach is built on the optimality of approximate message-passing algorithms among first-order iterative methods, delineating the fundamental learnability limit across a broad spectrum of algorithms, including neural networks trained with gradient descent.
1 Introduction
A fundamental property of neural networks is their ability to learn features from data and adapt to relevant structures in high-dimensional noisy data. However, our mathematical understanding of this mechanism remains limited. A popular model for studying this question is multi-index models. Multi-index functions are a class of statistical models encoding the inductive bias that the relevant directions for prediction depend only on a low-dimensional subspace of the covariates ${\bf x}\in\mathbb{R}^{d}$:
$y=g({{\bm{{W}}}}{\bm{x}}),\qquad{{\bm{{W}}}}\in\mathbb{R}^{p\times d},$ (1)
where the mapping $g:\mathbb{R}^{p}\rightarrow\mathbb{R}$ can further be stochastic i.e. depend on independent noise $\xi\in\mathbb{R}^{p^{\prime}}$ for some finte $p^{\prime}$. They define a rich class of hypotheses (Li, 1991), containing many widely studied functions in the statistical learning and theoretical computer science literature. The simplest instance is the Linear model ($p\!=\!1$) given by a linear link function $g(z)\!=\!z$, or its noisy version $g(z,\xi)\!=\!z\!+\!\sqrt{\Delta}\xi$, with $\xi\!\sim\!{\cal N}(0,1)$. In the Single-index model ($p=1$), also known as generalised linear model [^1] More precisely, the terminology single-index model is employed in a context where both the direction ${\bm{w}}$ and link function $g$ are learned, while generalised linear model is employed when the link function is known and only the weights are learned. In this work, we use both interchangeability., $g(z)$ can be an arbitrary, possibility stochastic, link function, including some widely studied problems such as phase retrieval $g(z)\!=\!z^{2}$ and the perceptron $g(z)={\rm sign}(z)$. Some example of multi-index models include the Two-layer neural network ($p>1$), given by a link function $g(z_{1},\dots,z_{p})=\sum_{k=1}^{p}a_{k}\sigma(z_{k})$; Polynomial functions ($p>1$), given by any linear combination of products of $z_{1},\dots,z_{p}$, for example $g(z_{1},z_{2},z_{3})=1+z_{1}^{2}z_{2}^{2}-2z_{1}+z_{2}z_{3}^{3}$; and s-sparse parities ($p>1$) that can be embedded in a multi-index model by taking a link function $g(z_{1},\dots,z_{p})=\prod_{k\in I}{\rm sign}(z_{k})$ for any subset $I\subset[p]$ of size $|I|=s\leq p$. Training a multi-index model typically translates to a non-convex optimization problem. Several authors have thus used multi-index models as a test-bed for understanding the behavior of neural nets and gradient-descent in non-convex, high-dimensional contexts, e.g. (Saad and Solla, 1995; 1996; Arous et al., 2021; Abbe et al., 2022; 2023; Veiga et al., 2022; Ba et al., 2022; Arnaboldi et al., 2023; Collins-Woodfin et al., 2023; Damian et al., 2023; Bietti et al., 2023; Moniri et al., 2023; Berthier et al., 2023).
To serve as a useful benchmark, it is necessary to have an idea of the fundamental limit of learnability in such models. This translates to the question of how many observations from model Equation 1 are required to obtain a better-than-random prediction within a class of algorithms, also known as weak learnability. This can be studied both statistically (within the class of all, including exponential time algorithms) or computationally (restricted to a particular computational class, such as first-order algorithms). In the single-index case ($p=1$), weak learnability has been heavily studied under probabilistic assumptions for the weights and data distribution (e.g. i.i.d. Gaussian or uniformly in the sphere). The statistical threshold for learnability has been characterised by Barbier et al. (2019) when the covariate dimension $d$ is large. Optimal computational thresholds for the class of first-order iterative algorithms were derived in (Mondelli and Montanari, 2019; Luo et al., 2019; Celentano et al., 2020; Maillard et al., 2020) in the same regime. Similar results were also proven for other computational models: Damian et al. (2022) provided a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. Damian et al. (2024) provided results for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$, hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent.
Aside from the particular case of committee machines (Aubin et al., 2019; Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), results for general multi-index models $p>1$ are scarce, as they crucially depend on the way different directions are coupled by the link function. The goal of the present work is precisely to close this gap for the class of Gaussian multi-index models in the proportional, high-dimensional regime, and to provide a classification of how hard it is to learn feature directions from data in multi-index models. More precisely, our main contributions are:
  • We analyse the computational limits of weak learnability (as defined in Def. 2) for Gaussian multi-index models in the class of first-order methods, in the high-dimensional limit when $d,n\to\infty$ at fixed ratio $\alpha=\nicefrac{{n}}{{d}}$ and constant number of indices $p=\Theta(1)$. Our analysis leverages the optimality of Bayesian approximate message passing algorithms (AMP) (Donoho et al., 2009; Rangan, 2011) among first-order iterative methods (Zdeborová and Krzakala, 2016; Celentano et al., 2020; Montanari and Wu, 2022), allowing us to delineate the fundamental learnability limits across a wide range of algorithms, including neural networks trained with gradient descent.
  • We provide a classification of which directions are computationally trivial, easy or hard to learn with AMP from linearly many samples in multi-index models: trivial directions can be learned with a single AMP iteration for any $\alpha>0$. When no such direction exists, we show there might be easy directions allowing recovery from arbitrarily small initial overlap $\epsilon>0$ in $\mathcal{O}(\log\frac{1}{\epsilon})$ steps with at least $n\!=\!\alpha_{c}d$ samples, where $\alpha_{c}$ marks a computational transition. Heuristically, we expect recovery along easy directions from random initialization with $\alpha>\alpha_{c}$ in $\mathcal{O}(\log d)$ iterations. We further confirm this numerically through extensive simulations.  [^2] Note that $\alpha_{c}$ may or may not coincide with the statistically optimal threshold. In no way do we imply the existence or absence of a computational-to-statistical gap. We provide a sharp formula for the critical sample complexity $\alpha_{c}$ above which easy directions are weakly learnable. It is conjectured that no efficient iterative algorithm can succeed for $\alpha<\alpha_{c}$. The hardest multi-index problems are those where $\alpha_{c}$ diverges, and where none of the directions can be learned from $n\!=\!\mathcal{O}(d)$ samples.
  • We discuss how interactions among directions can lead to hierarchical learning phenomena. Interestingly, even hard directions can be learned when coupled to easier ones through a hierarchical process, which is reminiscent of the staircase phenomenon for one-pass SGD (Abbe et al., 2023). However, the class of functions that are weakly learnable hierarchically with AMP differs from SGD and is, in particular, considerably larger. We characterise this class of Hierarchical Iterative Denoising Staircase functions and argue it is the relevant computational complexity class for learning multi-index models.
Further related work —
With its origins on the classical projection pursuit method (Friedman and Tukey, 1974; Friedman and Stuetzle, 1981), there is an extensive literature dedicated to designing and analysing efficient algorithms to train multi-index models models, such as Isotronic Regression for single-index (Brillinger, 1982; Kalai and Sastry, 2009; Kakade et al., 2011) and Sliced Inverse Regression in the multi-index case (Dalalyan et al., 2008; Yuan, 2011; Fornasier et al., 2012; Babichev and Bach, 2018).
The case $p=1$ has seen a lot of interest recently. In terms of gradient descent, Arous et al. (2021) has shown that if the link function $g$ is known, one-pass SGD achieves weak recovery in $n=\Theta(d^{\ell-1})$ steps, where $\ell$, known as the information exponent, is the first non-zero Hermite coefficient of the link function $g$. In the non-parametric setting —where $g$ is unknown— Berthier et al. (2023) has shown that a large-width two-layer neural network trained under one-pass SGD can learn an ”easy” single-index target ($\ell=1$) in $n=\Theta(d)$ steps. This is to be contrasted with full-batch GD, which can achieve weak recovery in $\Theta(1)$ steps with sample complexity $n=\Theta(d)$ even for particular problems with $\ell>1$ (Dandi et al., 2024) (which is, indeed, the statistical optimal rate (Barbier et al., 2019; Mondelli and Montanari, 2019)). Similar results have also been proven for abstract computational models. Damian et al. (2022) has proven a lower bound $n\geq d^{{\rm max}(1,\nicefrac{{\ell}}{{2}})}$ under the Correlational Statistical Query (CSQ) model, comprising algorithms that take queries of the type $\mathbb{E}[y\varphi({\bm{x}})]$. More recently, Damian et al. (2024) has proven a similar result for the Statistical Query (SQ) model, which allows for more flexible queries of the type $\mathbb{E}[\varphi({\bm{x}},y)]$ and hence a lower sample complexity $n\geq d^{{\rm max}(1,\nicefrac{{\kappa}}{{2}})}$ with $\kappa\leq\ell$ defining the generative exponent. This turns out to be equivalent to the optimal computational weak recovery threshold of (Barbier et al., 2019; Mondelli and Montanari, 2019; Luo et al., 2019; Maillard et al., 2020).
While the situation is less understood, the multicase has also whitenessed a surge of recent interest. In particular, it has been used to understand the behavior of gradiend descent algorithm in neural networks. For instance, Abbe et al. (2022; 2023) showed that a certain class of staircase functions can be learned by large-width two-layer networks trained under one-pass SGD with sample complexity $n=\Theta(d)$. This is in stark contrast to (embedded) $s$-sparse parities, which require $n\geq d^{s-1}$ samples (Blum et al., 2003). Abbe et al. (2022; 2023) introduced the leap exponent, a direction-wise generalisation of the information exponent for multi-index models, and studied the class of staircase functions which can be efficiently learned with one-pass SGD. Bietti et al. (2023) showed that under a particular gradient flow scheme preserving weight orthogonality (a.k.a. Stiefel gradient flow), training follow a saddle-to-saddle dynamics, with the characteristic time required to escape a saddle given by the leap exponent. Interestingly, it was then shown that the limit discovered in these set of works could be bypassed by slightly different algorithm, for instance by smoothing the landscapes Damian et al. (2023) or reusing batches multi-time Dandi et al. (2024). The latter paper, in particular, showing that gradient descent could learn efficiently a larger class of multi-index models that previously believed to be possible. These findings highlight the need of a strict understanding of the limit learnability of these models.
Our approach to establish the limit of computational learnability is based on the study of the approximate message passing algorithm (AMP). Originating from the cavity method in physics Mézard (1989); Kabashima (2008), AMP Donoho et al. (2009), and its generalized version GAMP Rangan (2011) are powerful iterative algorithm to study these high-dimensional setting. These algorithm are widely believed to be optimal between all polynomial algorithms for such high-dimensional problems Zdeborová and Krzakala (2016); Deshpande and Montanari (2015); Bandeira et al. (2018); Aubin et al. (2019); Bandeira et al. (2022). In fact, they are provably optimal among all iterative first-order algorithms Celentano et al. (2020); Montanari and Wu (2022), a very large class of methods that include gradient descent.
While these algorithm were studied in great detail for the single index models, see e.g. Barbier et al. (2019); Aubin et al. (2020b; a), and are at the roots of the spectral method that underline the learnability phase transition in this case Mondelli and Montanari (2019); Maillard et al. (2020), much less is known in the multi-index case, with the exception of Aubin et al. (2019), who showed that particular instances of multi-index models known as committee machines can be learned with $n=\mathcal{O}(d)$ samples.
Different from our approach are (Diakonikolas et al., 2020; Goel et al., 2020; Chen et al., 2022), who focused also on peculiar form of the multi-index models (using combination of Relu) and derived worst case bound (in terms of the hardest possible function). Instead, we focus on the typical-case learnability of a given, explicit, multi-index model.
Notation
[…]
2 Settings and definitions
As motivated in the introduction, our main focus in this work will be to study subspace identifiability in the class of Gaussian multi-index models.
Definition 1 (Gaussian multi-index models).
Given a covariate ${\bm{x}}\sim\mathcal{N}(0,\nicefrac{{1}}{{d}}{\bm{I}}_{d})$, we define the class of Gaussian multi-index models as output mappings of the type:
$y=g({{\bm{{W}}}}^{\star}{\bm{x}})$ (2)
where $g:\mathbb{R}^{p}\to\mathbb{R}$ denotes the link function and ${{\bm{{W}}}}^{\star}\in\mathbb{R}^{p\times d}$ is a weight matrix with i.i.d. rows ${\bm{w}}^{\star}_{k}\sim\mathcal{N}(0,{\bm{I}}_{d})$. Note we allow the map $g:\mathbb{R}^{p}\to\mathbb{R}$ to be stochastic i.e $g(\cdot)=\tilde{g}(\cdot,{\bm{\xi}})$ for random variable ${\bm{\xi}}\in\mathbb{R}^{p}$ independent of $\mathbf{x}$ having a fixed dimension-independent law such that $\mathbb{E}\left[\norm{\xi}_{2}^{2}\right]<\infty$. [^3] Technically, our results hold for ${\bm{\xi}}\in\mathbb{R}^{p^{\prime}}$ with $p^{\prime}=\mathcal{O}(1)$, for example the case of a degenerated noise. But for simplicity of exposition we focus on the $p^{\prime}=p$ case.
Given $n$ i.i.d. samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$ drawn as per Definition 1, we are interested in investigating the computational bottlenecks of estimating ${{\bm{{W}}}}^{\star}$ from the samples $({\bm{x}}_{i},y_{i})_{i\in[n]}$. Note that reconstructing ${{\bm{{W}}}}^{\star}$ or a permutation of its rows is equivalent from the perspective of the likelihood Equation 2. Therefore, in this work, we will be interested in weak subspace learnability, which corresponds to obtaining an estimation of the subspace spanned by ${{\bm{{W}}}}^{\star}$ better than a random estimator. This can be defined in an invariant way as follows
Definition 2 (Weak subspace recovery).
Let $V^{\star}\subset\mathbb{R}^{p}$ denote a subspace spanned by vectors representing components along ${{\bm{{W}}}}^{\star}$ such that each $\mathbf{v}\in V^{\star}$ maps to a vector $\mathbf{v}_{d}$ in $\mathbb{R}^{d}$ through the map $\mathbf{v}_{d}=({{\bm{{W}}}}^{\star})^{\top}\mathbf{v}$. Given an estimator $\hat{{{\bm{{W}}}}}\in\mathbb{R}^{p\times d}$ of ${{\bm{{W}}}}^{\star}$ with $\norm{\hat{{{\bm{{W}}}}}_{i}}^{2}_{F}=\Theta(d)$, we have weak recovery of a $V^{\star}$ if:
$\underset{{\bm{v}}\in\mathbb{R}^{p},\norm{{\bm{v}}}=1}{\inf}\norm{\frac{\hat{{{\bm{{W}}}}}({{\bm{{W}}}}^{\star})^{\top}{\bm{v}}}{d}}=\Theta(1).$ (3)
with high probability as $d\to\infty$.
Our main tool for characterising the computational bottlenecks in the Gaussian multi-index problem is an Approximate Message Passing (AMP) Algorithm 1 tailored to our problem. Note that Algorithm 1 only involves matrix multiplication by ${\bm{X}}\in\mathbb{R}^{n\times d}$ (and its transpose) and the application of a problem dependent function ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\in\mathbb{R}^{p}$ (defined later in Equation 4). Therefore, it belongs to the class of first-order algorithms with linear running time of $\mathcal{O}(nd)$. Intuitively, AMP estimates the mean ${\bm{\omega}}$ and variance ${\bm{V}}$ of ${{\bm{{W}}}}^{*}{\bm{x}}$ and the mean $\hat{{{\bm{{W}}}}}$ and covariance $\hat{{\bm{C}}}$ of ${{\bm{{W}}}}^{*}$ through a fixed point iteration described in Algorithm 1. The key property of AMP is that for a well-chosen ${\bm{g}}_{\rm out}$, it is provably optimal within the class of first-order methods (Celentano et al., 2020). This means that establishing learnability for the optimal AMP Algorithm 1 implies a computational lower bound in the class of first-order methods, which includes popular machine learning algorithms such as gradient descent (SG). For the Gaussian multi-index estimation problem 1, the optimal ${\bm{g}}_{\rm out}$ is simply given by the optimal denoiser of an effective $p$-dimensional problem $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
[…]
where the conditional expectation is defined through the output channel $Y=g({\bm{V}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}+{\bm{\omega}})$. For stochastic mappings $g(\cdot)=g(\cdot,\xi)$, the expectation is w.r.t both $Z,\xi$ conditioned on $Y$. Here ${\bm{\omega}}$ can be interpreted as an estimate of the pre-activations.
Our results rely on the following assumption on the above denoiser, translating to an indirect assumption on $g,\xi$:
Assumption 1.
${\bm{g}}_{\rm out}:\mathbb{R}^{p^{2}+p+1}\rightarrow\mathbb{R}^{p}\in\mathcal{C}^{2}$. Furthermore, the mapping $\hat{g}_{\rm out}:\mathbb{R}^{p+{p^{\prime}}}\rightarrow\mathbb{R}^{p}$ defined by ${\hat{g}}_{\rm out}(\mathbf{h},\xi)={\bm{g}}_{\rm out}(\tilde{g}((\mathbf{h},\xi),{\bm{\omega}},{\bm{V}})$ is pseudo-Lipschitz of finite-order for all ${\bm{\omega}},{\bm{V}}\in\mathbb{R}^{p},\mathbb{R}^{p\times p}$. Here, $\xi\in\mathbb{R}^{p^{\prime}}$ denotes independent noise satisfying $\mathbb{E}\left[\norm{\xi}^{2}\right]<\infty$ as per Definition 1.
The above denoiser can also be derived as an exact high-dimensional approximation of the Belief Propagation (BP) algorithm for the estimation of the marginals of the Bayesian posterior distribution:
$p({{\bm{{W}}}}|{\bm{X}},{\bm{y}})\propto\prod\limits_{i=1}^{n}\delta(y_{i}-g({{\bm{{W}}}}{\bm{x}}_{i}))\prod_{k=1}^{p}\mathcal{N}({\bm{w}}_{k}|0,{\bm{I}}_{d}).$ (5)
Algorithm 1 was derived for Gaussian multi-index models by Aubin et al. (2019), although it has been analysed only for the particular link function $g({\bm{z}})={\rm sign}(\sum_{k\in[p]}{\rm sign}(z_{k}))$, known as the committee machine. Our goal in this work is to leverage Algorithm 1 in order to provide a sharp classification of which link functions $g$ are computationally challenging to learn.
Algorithm 1 Multi-index AMP
  Input: Data ${\bm{X}}\in\mathbb{R}^{n\times d}$, ${\bm{y}}\in\mathbb{R}^{n}$
  Initialize $\hat{{{\bm{{W}}}}}^{t=0}\in\mathbb{R}^{p\times d}$, $\hat{{\bm{C}}}_{k}^{t=0}\in\mathcal{S}_{p}^{+}$ for $k\in[d]$, ${\bm{g}}^{t=0}\in\mathbb{R}^{n\times p}$.
  for $t\leq T$ do
     /* Update likelihood mean and variance
     ${\bm{V}}_{i}^{t}=\sum_{k=1}^{d}X^{2}_{ik}\hat{{\bm{C}}}_{k}\in\mathbb{R}^{p\times p}$;  $\mathbf{\omega}_{i}^{t}=\sum_{k=1}^{d}X_{ik}\hat{{{\bm{{W}}}}}_{k}^{t}-{\bm{V}}_{i}^{t}{\bm{g}}_{i}^{t-1}\in\mathbb{R}^{p}$, $i\in[n]$;
     ${\bm{g}}_{i}^{t}=g_{\rm{\rm out}}(y_{i},{\bm{\omega}}_{i}^{t},{\bm{V}}_{i}^{t})\in\mathbb{R}^{p}$ ;  $\partial{\bm{g}}_{i}^{t}=\partial_{{\bm{\omega}}}g_{\rm{\rm out}}(y_{i},{\bm{\omega}}_{i}^{t},{\bm{V}}_{i}^{t})\in\mathbb{R}^{p\times p}$ ; $i\in[n]$
     /* Update prior first and second moments
     ${\bm{A}}_{k}^{t}=-\sum_{i=1}^{n}X_{ik}^{2}\partial{\bm{g}}^{t}_{i}\in\mathbb{R}^{p\times p}$ ;   $\bm{b}_{k}^{t}=\sum_{i\in[n]}X_{ik}{\bm{g}}_{i}^{t}+{\bm{A}}_{k}^{t}\hat{{{\bm{{W}}}}}_{k}^{t}$;  $k\in[d]$
     $\hat{{{\bm{{W}}}}}_{k}^{t+1}=({\bm{I}}_{p}+{\bm{A}}_{k}^{t})^{-1}\bm{b}_{k}^{t}\in\mathbb{R}^{p}$ ;  $\hat{{\bm{C}}}_{k}^{t+1}=({\bm{I}}_{p}+{\bm{A}}_{k}^{t})^{-1}\in\mathbb{R}^{p\times p}$,  $k\in[d]$
  end for
  Return: Estimators $\hat{{{\bm{{W}}}}}_{\rm amp}\in\mathbb{R}^{p\times d},\hat{{\bm{C}}}_{\rm amp}\in\mathbb{R}^{d\times p\times p}$
Our analysis is based on two remarkable properties that make AMP a particularly useful tool for studying high-dimensional estimation. The first property is that for any $t=\Theta(1)$, in the high-dimensional limit $d\to\infty$ the performance of AMP can be tracked without actually running the algorithm.
This result, known as state evolution, makes AMP mathematically tractable in high-dimensions (Bayati and Montanari, 2011). The second property is its optimality with respect to Bayesian estimation. We discuss this further in Appendix A.
For multi-index models, the state evolution equations were derived by Aubin et al. (2020b) and rigorously proven by Gerbelot and Berthier (2023). The state evolution equations, in particular, provide an exact characterization of the asymptotic overlaps and prediction error.
Lemma 2.1 (State evolution (Aubin et al., 2019; Gerbelot and Berthier, 2023)).
Let $({\bm{x}}_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Consider running Algorithm 1 from initial condition $\hat{{{\bm{{W}}}}}^{0}\in\mathbb{R}^{p\times d}$, such that the initial overlap $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ converges in probability to a limit ${{\bm{M}}}^{0}$ as $d\rightarrow\infty$. Denote by $\hat{{{\bm{{W}}}}}^{t}$ the resulting estimator at time $0\leq t\leq T$. Then, in the high-dimensional limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, constant $p$ and any finite time $T$, the limiting overlaps satisfy:
[…]
with ${{\bm{M}}}^{t}$ satisfying the following state evolution equations from initial condition ${{\bm{M}}}^{0}$:
${{\bm{M}}}^{t+1}\!$ $=\!F({{\bm{M}}}^{t})\!\coloneqq{G}\!\left(\!\alpha\,\mathbb{E}\Big{[}{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}\!\Big{]}\right).$ (7)
where $G({\bm{X}})=({\bm{I}}_{p}+{\bm{X}})^{-1}{\bm{X}}$ and the expectation is taken over the following effective process
$Y^{t}=g\left(({\bm{I}}_{p}-{{\bm{M}}}^{t})^{\nicefrac{{1}}{{2}}}{\bm{\xi}}+{{{\bm{M}}}^{t}}^{\nicefrac{{1}}{{2}}}{\bm{Z}}\right),\qquad{\bm{Z}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ (8)
with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})$ independently from ${\bm{Z}}$. Furthermore, ${{\bm{M}}}^{t}$ is constrained to lie in $\mathcal{S}^{+}_{p}$ due to Nishimori-identity (section G.3 in (Aubin et al., 2019)).
As a consequence, the asymptotic mean-squared error on the label prediction is given by:
${\mathbb{E}_{\bm{X},\bm{y}}}\mathbb{E}_{\bm{x},y}\left[\left(y-g\left(\hat{\bm{W}}^{t}(\bm{X},\bm{y})\bm{x}\right)\right)^{2}\right]=\mathbb{E}[(Y^{t}-g({\bm{Z}}))^{2}],$ $0\leq t\leq T$ (9)
where the expectation is taken over the effective estimation process Equation 8.
3 The trivial subspace ($\alpha_{c}=0$)
The state evolution Equation 7 maps the problem of characterising the computational bottlenecks of first-order methods for high-dimensional Gaussian multi-index models to the study of the deterministic, $p$-dimensional dynamical system ${{\bm{M}}}^{t+1}=F({{\bm{M}}}^{t})$. A starting point is identifying its fixed points and their basins of attraction. In the absence of any prior information on ${{\bm{{W}}}}^{\star}$ aside from its distribution, one cannot do better than taking $\hat{{\bm{w}}}_{k}^{t=0}\sim\mathcal{N}(0,{\bm{I}}_{d})$ with $k\in[p]$ independently at random from the prior. With high-probability, at initialisation, the elements of the overlap matrix $\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{0}{{{\bm{{W}}}}^{\star}}^{\top}$ are $\Theta(d^{-\nicefrac{{1}}{{2}}})$ element-wise. The asymptotic overlap for an uninformed initial condition is thus identically zero ${{\bm{M}}}^{0}=0$, a null-rank matrix. If ${{\bm{M}}}^{0}=0$ is not a fixed point, then ${{\bm{M}}}^{1}\succ 0$, implying the weak recovery of a subspace of dimension $k={\rm rank}({{\bm{M}}}^{1})>0$ with just a single step of AMP.
Lemma 3.1 (Existence of uninformed fixed point).
${{\bm{M}}}=0\in\mathbb{R}^{p\times p}$ is a fixed point of Equation 7 if and only if the following condition holds almost surely over $Y$:
${\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})=\mathbb{E}[{\bm{Z}}|Y]=\mathbf{0},$ (10)
This implies that as long as the conditional expectation above is not zero almost surely, AMP weakly learns a non-empty subspace immediately in the first iteration for any number of samples $n=\Theta(d)$. For this reason, we refer to this subspace as a “trivial subspace”:
Definition 3 (Trivial subspace).
We define $H^{\star}_{T}\subseteq\mathbb{R}^{p}$ as the subspace spanned by $\mathbf{v}\in\mathbb{R}^{p}$ such that the following holds almost surely over $Y=g({\bm{Z}})$ with ${\bm{Z}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$:
${\bm{g}}_{\rm out}(Y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}[\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle|y=Y]=0,$ (11)
where the expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1. Equation 8. The trivial subspace $T^{\star}$ is defined as the orthogonal complement of $H^{\star}$.
Note that for single-index models ($p=1$), the condition in Equation 10 reduces to the one derived by (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020). Interestingly, this is exactly the same condition appearing in Damian et al. (2024) for weak learnability of single-index models in the SQ model; see eq. (3) therein. To make Lemma 3.1 concrete, let us look at a few examples. A detailed derivation of these examples is discussed in the Appendix C.
  1. 1.
    For single-index models ($p=1$), $T^{*}$ is one dimensional if and only if $g$ is non-even, e.g. $g(z)={\rm He}_{3}(z)$. This follows from requiring that $g_{\rm out}(y,0,1)\neq 0$ for at least one value of $y$. In particular, on any open interval where $g_{\rm out}$ is invertible we have $g_{\rm out}=g^{-1}$.
  2. 2.
    For a linear multi-index model, $g({\bm{z}})=\sum_{i=1}^{p}z_{i}$, $T^{\star}$ is spanned by $\mathbf{1}_{p}\in\mathbb{R}^{p}$ (all-one vector).
  3. 3.
    For a committee $g({\bm{z}})\!=\!\sum_{i=1}^{p}{\rm sign}(z_{i})$, the trivial subspace $T^{\star}$ is again 1-d, spanned by $\mathbf{1}\!\in\!\mathbb{R}^{p}$.
  4. 4.
    For monomials $g({\bm{z}})=z_{1}\dots z_{p}$, the trivial subspace $T^{\star}$ is non-empty if and only if $p=1$.
  5. 5.
    For staircase functions (Abbe et al., 2023):
    $g({\bm{z}})=z_{1}+z_{1}z_{2}+z_{1}z_{2}z_{3}+\dots+z_{1}\dots z_{p}$ (12)
    The trivial subspace is $T^{\star}=\mathbb{R}^{p}$ and is spanned by the canonical basis. In other words, AMP learns all the directions with a single step for any $\alpha>0$.
Interestingly, Lemma 3.1 can also be related to computational models based on queries, such as SQ learning (Kearns, 1998): the denoiser ${\bm{g}}_{\rm out}$ can indeed be interpreted as a non-linear transformation on the labels $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$. From this perspective, the statement on the condition for the existence of a non-empty trivial subspace Equation 11 translates to the following condition:
$\mathbb{E}\left[{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=\mathbb{E}\left[\mathbb{E}[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|Y=y]^{2}\right]\neq 0,$ (13)
where ${\bm{v}}\in T^{\star}$. Indeed, the left-hand side can be interpreted as a statistical query of the type $\mathbb{E}[\varphi(y)\psi(x)]$ with label pre-processing $\varphi={\bm{g}}_{\rm out}$. The fact that this linear correlation in the transformed labels is non-vanishing implies that one can weakly recover $\mathbf{v}$ through a tailored spectral method (Mondelli and Montanari, 2019). In fact the denoiser ${\bm{g}}_{\rm out}$ is the optimal such transformation in the sense that when ${\bm{g}}_{\rm out}$ fails to obtain a linear correlation along $\mathbf{v}$, i.e when $\mathbf{v}\in H^{\star}$, then no transformation can:
Lemma 3.2.
For any $\mathbf{v}\in H^{\star}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle\right]=0$ (14)
Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem).
Proof.
The above is a direct consequence of the Tower law of expectation. Specifically, we have:
$\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right].$
The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$. ∎
This implies that when $T^{\star}$ is empty, running one-pass SGD after applying any transformation $\mathcal{T}$ fails to obtain weak recovery with $\mathcal{O}(d)$ samples, as the absence of a first-order Hermite coefficient implies a sample complexity of $\mathcal{O}(d\log d)$ as shown in Theorems 1.3 and 1.4 in Arous et al. (2021). The situation can be very different when reusing batches: Dandi et al. (2024) showed that full batch GD implicitly applies a transformation $\mathcal{T}$ to the labels, in a mechanism similar to the one discussed above. However, this transformation may not be optimal. For instance, while AMP learns the trivial subspaces of both $g(z)={\rm He}_{3}(z)$ and staircase functions (12) in a single-step, GD requires 2-steps for the first and $p$ steps for the latter (Dandi et al., 2024).
4 Computational phase transitions ($\alpha_{c}>0$)
What happens when the trivial subspace $T^{\star}$ from Definition 3 is empty? In this section, we discuss precisely this case, showing that for some link functions $g$ there exists a critical sample complexity threshold $\alpha_{c}>0$ above which some directions become learnable by iterating AMP. To contrast with the trivial subspace, we refer to these as the easy directions. While the trivial subspace is characterised by the existence of the fixed point ${{\bm{M}}}=0$, the easy subspace will be determined by its stability, which crucially depends on the sample complexity $\alpha=n/d$. The stability of ${{\bm{M}}}=\mathbf{0}$ can be obtained by linearising the state evolution (7) around $\delta{{\bm{M}}}$ with $\norm{\delta{{\bm{M}}}}\approx 0$
$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\mathcal{O}(\norm{\delta{{\bm{M}}}}^{2})$ (15)
where $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\mathbb{E}_{Y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,0,{\bm{I}}_{p})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,0,{\bm{I}}_{p})^{\top}\right]\,,$ (16)
This can be summarized as follows:
Lemma 4.1.
(Stability of the uninformed fixed point). If ${{\bm{M}}}=\mathbf{0}\in\mathbb{R}^{p\times p}$ is a fixed point, then it is an unstable fixed point of Equation 7 if and only if $n>\alpha_{c}d$, where the critical sample complexity $\alpha_{c}$ is:
$\frac{1}{\alpha_{c}}=\sup_{{{\bm{M}}}\in\mathcal{S}^{+}_{p}}\norm{\mathcal{F}({{\bm{M}}})}_{F},$ (17)
with $\|\cdot\|_{F}$ the Frobenius norm, and the extremum being achieved at a unique extremizer ${{\bm{M}}}^{\star}\in\mathcal{S}^{+}_{p}$. Moreover, if $\norm{\mathcal{F}({{\bm{M}}})}_{F}=0$, then ${{\bm{M}}}=\mathbf{0}$ is a stable fixed point for any $n=\Theta(d)$.
The above result follows from a generalization of the Perron-Frobenius theorem to the cone $\mathcal{S}^{+}_{p}$; see Section B.2 for a proof. A consequence of this result is that for $\alpha<\alpha_{c}$, even with infinitesimally small informed initialisation AMP will converge back to the trivial fixed point and fail to recover any subspace, while for $\alpha>\alpha_{c}$, it develops a growing overlap along a non-empty subspace starting with arbitrarily small initialization. This marks the onset of a computational phase transition for weak learnability in the first-order methods class. Note that the computational weak learnability threshold $\alpha_{c}$ can be equal to the information-theoretical one (as, for instance, in ”real” phase retrieval (Mondelli and Montanari, 2019; Maillard et al., 2020)), and by no means we imply the existence (or lack of thereof) of a computational-to-statistical gap (Zdeborová and Krzakala, 2016). The easy subspace spanned by the directions that can be learned with $n>\alpha_{c}d$ samples can be characterised as follows:
Definition 4 (Easy subspace $E^{\star}$).
Let $H^{*}_{E}$ be the subspace spanned by directions $\mathbf{v}\in\mathbb{R}^{p}$ such that:
$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=0,{\bm{V}}={\bm{I}}_{p})\mathbf{v}=0,$ (18)
almost surely over $Y$. We define the easy subspace $E^{\star}$ as the orthogonal complement of $H^{\star}_{E}$.
Lemma 4.1 implies that for $\alpha>\alpha_{c}$ AMP Algorithm 1 initialised from small but non-zero ${{\bm{M}}}^{0}$ will learn some directions in the easy subspace $E^{\star}$. Surprisingly, we can show that for $\alpha>\alpha_{c}$ AMP Algorithm 1 spans the full easy subspace $E^{\star}$ starting from an arbitrarily small but extensive initial correlation.
Theorem 4.2.
Let ${{\bm{M}}}_{d}^{t}\coloneqq\nicefrac{{1}}{{d}}\hat{{{\bm{{W}}}}}^{t}{{{\bm{{W}}}}^{\star}}^{\top}$ denote the model-target overlap matrix at any finite time $t$. Suppose that $T^{\star}={0}$ and the AMP algorithm 1 is initialised at $\hat{{{\bm{{W}}}}}^{0}$ such that ${{\bm{M}}}_{d}^{0}=\epsilon{\bm{A}}$, with $\norm{{\bm{A}}}_{F}=1$ and ${\bm{A}}$ being full-rank. Then,under Assumption 1, with high probability as $d\rightarrow\infty$, for $\alpha\geq\alpha_{c}$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}^{\star}$ for $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{\star}$ is as defined in Lemma 4.1. For $\alpha<\alpha_{c}$ however, ${{\bm{M}}}^{t}_{d}=0$ is asympotically stable i.e. there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{d}}\rightarrow 0$ as $t\rightarrow\infty$. Furthermore, there exists an $\alpha>0$ and a $\delta>0$ such that ${{\bm{M}}}_{d}^{t}\succ\delta M_{E^{*}}$ in $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$ iterations, where $M_{E^{\star}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}$.
Note that in the above the initial overlap scale $\epsilon>0$ can be arbitrarily small but constant with respect to $d$. Heuristically, $\epsilon$ is often identified with $\mathcal{O}(\nicefrac{{1}}{{\sqrt{d}}})$, the scale of a random initial condition. Although this heuristic is not covered by Theorem 4.2, we have numerically checked it is consistent with the extrapolation of the theory, which suggests that for $\alpha>\alpha_{c}$ easy subspaces can be weakly recovered with $\mathcal{O}(\log d)$ steps, see Figure 2.
As for Lemma 3.1, the expression for the weak recovery threshold Equation 17 generalises the single-index expression from (Mondelli and Montanari, 2019; Barbier et al., 2019; Maillard et al., 2020). Indeed, for $p=1$, we have $\partial_{\omega}g_{\rm out}(y,0,1)=\mathbb{E}[{\rm He}_{2}(Z)|Y=y]$, and therefore
[…]
which is exactly eq. (11) in Barbier et al. (2019). Optimality of AMP then implies that no first-order method can achieve non-vanishing correlation with ${\bm{w}}^{\star}$ when $n<\alpha_{c}d$ as $d\to\infty$.
Figure 1: Numerical illustration of the weak learnability phase transition for the 2-sparse parity $g(z_{1},z_{2})\!=\!{\rm sign}(z_{1}z_{2})$ that has a phase transition at $\alpha_{c}(2)=\pi^{2}/4$. The overlap shows how well the directions $z_{1}$ and $z_{2}$ are recovered. Given the permutation symmetry in (20), we show here and in all the subsequent figures the optimal permutation of the overlap matrix elements reached by AMP. The solid black line is the prediction from the theory. Crosses are averages over $72$ runs of AMP Algorithm 1 with $d\!=\!500$.
We now illustrate Lemma 4.1 in a few examples of interest.
  1. 1.
    The monomial $g({\bm{z}})=z_{1}\dots z_{p}$ with $p>1$ can always be learned with $\alpha>\alpha_{c}(p)$ large enough (Chen and Meka, 2020). For instance, we have $\alpha_{c}(2)\approx 0.5937$, $\alpha_{c}(3)\approx 3.725$ and $\alpha_{c}(4)\approx 4.912$. In Appendix C we derive an analytical formula for $\alpha_{c}(p)$ for arbitrary $p$, and show that $\alpha_{c}(p)\sim p^{1.2}$ for large $p$.
  2. 2.
    The embedding of the sparse parity functions:
    $g({\bm{z}})=\prod_{k=1}^{p}{\rm sign}(z_{k})\,.$ (20)
    As this is invariant under permutations of the indices, Lemma 4.1 implies the existence of a computational phase transition. In Section C.3 we compute analytically the critical value $\alpha_{c}(p)$. For $p=1$ this problem is equivalent to phase retrieval, for which we have $\alpha_{c}(1)=\nicefrac{{1}}{{2}}$. For $p=2$ we show in the appendix that $\alpha_{c}(2)=\nicefrac{{\pi^{2}}}{{4}}$, while $\alpha_{c}(p)=+\infty$ for $p\geq 3$. Figure 1 illustrates this discussion for the 2-sparse parity problem, comparing theoretical prediction Equation 7 with finite-size runs of Algorithm 1.
Figure 2: Hierarchical weak learnability for the staircase function $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$. (Left): Overlaps with the first direction $|M_{11}|$ ( blue), and with the second and third one $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ ( red) as a function of the sample complexity $\alpha=\nicefrac{{n}}{{d}}$, with solid lines denoting state evolution curves Equation 7, and crosses/dots finite-size runs of AMP Algorithm 1 with $d=500$ and averaged over $72$ seeds. All other overlaps are zero (black). The two black dots indicate the critical thresholds at $\alpha_{1}\approx 0.575$ and $\alpha_{2}=\pi^{2}/4$. (Right) Corresponding generalization error as a function of the sample complexity. Details on the numerical implementation are discussed in Appendix D.
The sparse-parity function is a classic example of a computationally hard function in theoretical computer science (Blum et al., 1994; 2003). [^4] Strictly speaking, the noiseless parity is efficiently learnable by Gaussian elimination (Blum et al., 1994). Here, we consider computational efficiency in a noise-tolerant sense. Therefore, it should come as no surprise that $\alpha_{c}(p)=\infty$ for $p\geq 3$, providing a concrete example of what a hard subspace (i.e. that cannot be learned with $n=\Theta(d)$ samples) looks like for AMP. Similar to the discussion of trivial directions in Lemma 3.2, we can invoke optimality of AMP to translate our result on the optimal denoiser $y\mapsto{\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$ to a statement for queries on general label pre-processing transformations:
Lemma 4.3.
For any $\mathbf{v}\in H^{\star}_{E}$ and any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$:
$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0$ (21)
Note that the expectation is with respect to the distribution of the labels $y$ (not the effective problem from Equation 8).
For an AMP-hard function (that is, when $\alpha_{c}=\infty$), the above Lemma implies that, even after applying any transformation to the output (as allowed in SQ) online-SGD would require at least $\Omega(d^{2})$ data, since online-SGD with squared loss requires $\Omega(d^{\kappa-1})$ samples/iterations for weak-recovery of a target function when the first $\kappa$ Hermite-coefficients are 00 (Theorems 1.3 and 1.4 in Arous et al. (2021)). Furthermore, for single-index models satisfying the condition in Lemma 4.3 (i.e $\alpha_{c}=\infty$), the SQ lower-bound predicts a sample complexity requirement of at-least $\Omega(d^{3/2})$ (Damian et al., 2024). Therefore, learning such functions with $\mathcal{O}(d)$ samples is conjectured to be algorithmically-hard.
5 Hierarchical Iterative Denoising
Figure 3: Trajectories of a single finite-size run of AMP with $d=500$ at $\alpha=4$ for $g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$. (Left) Evolution of the overlaps. We display $M_{11}$ in blue, $\nicefrac{{1}}{{2}}(M_{22}+M_{33})$ in red, and the off-diagonal overlaps in black. (Right) Evolution of the generalisation error.
In Section 3 and 4 we have provided a full classification of what class of multi-index models are trivial, easy or hard to learn for a first-order algorithm. This classification was built upon the notion of weak learnability, or the capacity of AMP to learn subspaces starting from a random initial condition. However, it says nothing about about what can be learned after a given subspace is learned, for instance, from possible side information about a particular subspace. We now address this question. Suppose that the estimator $\hat{{{\bm{{W}}}}}^{t}$ has developed an overlap along a subspace belonging to the span of ${{\bm{{W}}}}^{\star}$, resulting in a non-zero overlap ${{\bm{M}}}^{t}\succ 0$. From that point, the main difference with respect to the previous discussion is that the variable ${\bm{\omega}}$ in the linear operator $\mathcal{F}$ defined in Equation 16 becomes non-zero (since it is distributed as ${\bm{\omega}}=\sqrt{{{\bm{M}}}}{\bm{\xi}}$). Crucially, this changes the span of $\mathcal{F}({{\bm{M}}})$ and hence the stability condition in Lemma 4.1. In particular, as we show next, learning some directions might facilitate learning larger subspaces. This phenomenon is reminiscent of the specialisation transition in committee machines (Saad and Solla, 1996; Aubin et al., 2020b) and of the staircase phenomenon for one-pass SGD introduced in (Abbe et al., 2022). Below, we formalise this discussion:
Definition 5.
Let $U\in\mathbb{R}^{p}$ be a subspace of dimensions $k$. We define $H_{E}^{\star}(U)$ to be the subspace spanned by ${\bm{v}}\in U^{\perp}$ such that:
${\bm{v}}^{\top}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}}_{U}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U})=0,$ (22)
almost surely over ${\bm{\xi}}\sim\mathcal{N}(\mathbf{0},{\bm{I}}_{p})$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$. We define the “trivially-denoising-coupled” subspace $T^{\star}_{U}$ for $U$ as the orthogonal complement of $H_{E}^{\star}(U)$.
Analogously, let $H^{\star}_{E}(U)$ be the subspace spanned by directions ${\bm{v}}\in U^{\perp}$ such that:
${\bm{v}}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{{{\bm{M}}}}_{U}){\bm{v}}=0,$ (23)
almost surely over ${\bm{\xi}}$ and $Y$ for any ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ such that $\operatorname{span}({{\bm{M}}}_{U})=U$.
When ${{\bm{M}}}_{U}$ is additionally a fixed point of $\mathcal{F}_{{{\bm{M}}}}$, we can linearise $\mathcal{F}_{{{\bm{M}}}}$ along the orthogonal complement of $U$. We define the easy-denoising-coupled subspace $E^{\star}_{U}$ for $U$ as the orthogonal complement of $H^{\star}_{E}(U)$. Next, suppose that ${{\bm{M}}}_{U}\in\mathcal{S}^{+}_{p}$ with $\operatorname{span}({{\bm{M}}}_{U})=U$ is further a fixed-point of $\mathcal{F}_{{{\bm{M}}}}$. Let $\mathcal{F}_{{{\bm{M}}}_{U}}$ denote the linearization of $F({{\bm{M}}})$ along the orthogonal complement $U^{\perp}$ at ${{\bm{M}}}={{\bm{M}}}_{U}$. We define the iterative denoising critical threshold $\alpha_{\text{hid}}({{\bm{M}}}_{U})$ at ${{\bm{M}}}={{\bm{M}}}_{U}$ as:
$\frac{1}{\alpha_{\text{hid}}({{\bm{M}}}_{U})}=\sup_{{{\bm{M}}}^{\perp}\in U^{\perp}}\norm{\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}})}_{F},$ (24)
We denote by ${{\bm{M}}}^{\star}_{U}$ the extremiser achieving the above supremum.
Figure 4: Specialisation as a staircase with $g\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$. The direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is first learned for any $\alpha>0$. The remaining ones are only learned at the specialisation transition occurring at $\alpha\!\approx\!4.3$. We take care of the symmetries as indicated in Appendix D. Crosses denote AMP runs with $d\!=\!100$ averaged over $72$ seeds. The overlap is shown as a function of the sample complexity. Because of the symmetry the elements of $M$ can take one of two values: one on the diagonal and one outside of it. We display $(M_{11}+M_{22}+M_{33})/3$ in blue and $(M_{12}+M_{13}+M_{23})/3$ in red.
The definitions above generalise the notions of trivial and easy subspaces in Definitions 3 and 4 conditionally on a subspace $U$ that has been previously learned, and characterise the directions whose recovery is enabled upon learning the subspace $U$. Concretely, upon developing an initial overlap along $U$, the directions in $T^{\star}_{U},E^{\star}_{U}$ can be recovered analogous to the recovery of $T^{\star},E^{\star}$ starting from random initialization. We formalise this below:
Theorem 5.1.
Let $U\in\mathbb{R}^{p}$ be a fixed subspace. Suppose the AMP algorithm 1 is initialised such that ${{\bm{M}}}_{d}^{0}={{\bm{M}}}_{U}+\epsilon{\bm{A}}$, where $\norm{{\bm{A}}}_{F}=1$ and ${{\bm{M}}}_{U}$ is a fixed point of $F({{\bm{M}}})$ in Equation 7 with $\operatorname{span}({{\bm{M}}}_{U})=U$. Then, under Assumption 1, the following holds with high probability as $d\rightarrow\infty$: If $\alpha\geq\alpha_{\text{hid}}({{\bm{M}}}_{U})$, $\exists\delta>0$ such that for any sufficiently small $\epsilon$ ,${{\bm{M}}}^{(t)}\succ\delta{{\bm{M}}}^{*}_{U}$ for some $t=\mathcal{O}(\log\nicefrac{{1}}{{\epsilon}})$, where ${{\bm{M}}}^{*}_{U}$ is defined as in Definition 5. For $\alpha<\alpha_{\text{hid}}({{\bm{M}}}_{U})$ however, there exists an $\epsilon^{\prime}$ such that for $\epsilon<\epsilon^{\prime}$, $\norm{{{\bm{M}}}^{t}_{\perp}}=0$ as $t\rightarrow\infty$, where ${{\bm{M}}}^{t}_{\perp}$ denotes the projection of ${{\bm{M}}}^{t}$ orthogonal to $U$. Furthermore, suppose that ${\bm{A}}$ is full-rank, then there exists an $\alpha>0$ and a $\delta>0$ such that is ${{\bm{M}}}_{d}^{t}\succ\delta{{\bm{M}}}_{E^{\star}_{U}}$, where ${{\bm{M}}}_{E^{\star}_{U}}\in\mathcal{S}^{+}_{p}$ spans $E^{\star}_{U}$
When $E^{\star}_{U}$ is non-empty for some subspace $U\subseteq\mathbb{R}^{p}$, we say that the target $y=g({{\bm{{W}}}}{\bm{x}})$ allows learning through Hierarchical Iterative Denoising. A concrete example of a function which displays the hierarchical iterative denoising staircase is a linear combination between hard parity function and an easy polynomial:
$g(z_{1},z_{2},z_{3})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ (25)
The sign part is a sparse parity with $p=3$, which cannot be learned with $n=\Theta(d)$ samples, but the quadratic polynomial $z_{1}^{2}$ component in the function allows weak-recovery of the first component i.e. $U={(1,0,0)}$ as long as $\alpha>1/2$. Hence, conditionally on $U$ the effective multi-index model becomes ${\rm sign}(z_{2}z_{3})$, which as discussed in Section 4 is an easy function. Figure 2 illustrates this: first, $z_{1}$ is learned at $\alpha_{1}\approx 0.575$. Then, for larger value when $\alpha>\alpha_{2}$, all directions are learned (see App. (C.4)). Knowing $z_{1}$ makes the hard 3-parity an easy 2-parity problem. We can see the evolution of the overlaps over time in Figure 3.
It is easy to construct a multi-index model where AMP will iterate over any number of such plateaus. For instance, consider the model:
$g({\bm{z}})=z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})+{\rm sign}(z_{3}z_{4}z_{5})+{\rm sign}(z_{5}z_{6}z_{7})+\ldots$ (26)
After the first plateau to learn $z_{1}$, there will be one for $z_{2}$, then for $z_{3}$, for $z_{4}$ and so on.
Another example is provided in Fig. 4 for the neural net $g(z_{1},z_{2},z_{3})\!=\!{\rm sign}(z_{1})\!+\!{\rm sign}(z_{2})\!+\!{\rm sign}(z_{3})$. At first only the one-dimensional direction spanned by $z_{1}\!+\!z_{2}\!+\!z_{3}$ is learned for any $\alpha>0$. This is simply a linear approximation of the function. Only at $\alpha\!\approx\!4.3$ does the network learns the other directions. Known as the the specialisation transition, this has been first discussed by Saad and Solla (1996); Aubin et al. (2020a), and provide a very early example of iterative sequential learning.
Finally, it is important to stress that hierarchical learnable functions are not equivalent to staircase functions (Abbe et al., 2022; 2023). For instance, consider the target:
$g({\bm{z}})=\rm He_{4}(z_{1})+{\rm sign}(z_{1}z_{2}z_{3}).$ (27)
Similar to Example 26, the above target is learned sequentially with $\mathcal{O}(d)$ samples by first recovering the direction corresponding to $z_{1}$. However, under the CSQ staircase classification, $g({\bm{z}})$ corresponds to leap $3$. The CSQ lower-bound therefore implies a sample complexity-requirement of $\Omega(d^{1.5})$ with online-SGD learning it in $\mathcal{O}(d^{2})$ samples (Abbe et al., 2023).
As we argued in Lemmas 3.2 and 4.3, we believe Hierarchical Iterative Denoising provides the right notion of computational complexity for these models, since it encompasses a larger class of natural algorithms such as batch gradient descent.
6 Conclusion and limitations
This work investigated the computational boundaries of learning high-dimensional Gaussian multi-index models with first-order iterative algorithms. As we hinted in the text (and provided some partial answers), a natural question is how close neural networks are to achieving this optimal performance. Two limitations of our work are: (a) our results hold asymptotically in $d\to\infty$; (b) We assume Gaussian covariates. For the former, numerical experiments support that the predictions are already accurate at fairly moderate sizes. For the latter, a number of universality results justify the scope of high-dimensional Gaussian covariates and extend AMP to more structured designs (Loureiro et al., 2021; Wang et al., 2022; Dudeja et al., 2023; Dandi et al., 2023).
Acknowledgements
The authors would like to thank Joan Bruna, Theodor Misiakiewicz, Luca Pesce, and Nati Srebro for stimulating discussions. This work was supported by the Swiss National Science Foundation under grants SNSF SMArtNet (grant number 212049) and SNSF OperaGOST (grant number 200390) and the Choose France - CNRS AI Rising Talents program.
Appendix / supplemental material
Appendix A Optimality of AMP
Recall that AMP Algorithm 1 is tailored to estimate the posterior marginals Equation 5. But how efficient is it? A priori, making this comparison requires sampling from the posterior distribution, which is computationally prohibitive in the high-dimensional limit of interest here. The second remarkable property of AMP is that its optimality with respect to the Bayesian posterior Equation 5 can be exactly characterised in the high-dimensional regime.
Lemma A.1 (Bayes-optimal correlation).
Let $(x_{i},y_{i})_{i\in[n]}$ denote $n$ i.i.d. samples from the multi-index model defined in 1. Denote by $\hat{{{\bm{{W}}}}}_{\rm bo}=\mathbb{E}[W|X,y]\in\mathbb{R}^{p\times d}$ the mean of the posterior marginals Equation 5. Then, under Assumption 1 in the high-dimensional asymptotic limit where $n,d\to\infty$ with fixed ratio $\alpha=\nicefrac{{n}}{{d}}$, the asymptotic correlation between the posterior mean and ${{\bm{{W}}}}^{\star}$:
${{\bm{M}}}^{\star}=\lim\limits_{d\to\infty}\mathbb{E}\left[\frac{1}{d}\hat{{{\bm{{W}}}}}_{\rm bo}{{{\bm{{W}}}}^{\star}}^{\top}\right]$ (28)
is the solution of the following $\sup\inf$ problem:
[…]
where $H_{Y}({{\bm{M}}})=\mathbb{E}_{{\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}}_{p})}[H_{Y}({\bm{m}}|{\bm{\xi}})]$, with $H_{Y}({{\bm{M}}}|{\bm{\xi}})$ the the conditional entropy of the effective $p$-dimensional estimation problem Equation 8.
Note that the state evolution Equation 7 is closely related to the $\sup\inf$ problem in Equation 29. Indeed, remarking that the update function $F$ in Equation 7 is precisely the gradient of the entropy $H_{Y}$ in Equation 8, one can show that state evolution is equivalent to gradient descent in the objective defined by Equation 29 [Aubin et al., 2019]. This non-trivial fact implies that whenever Equation 29 has a single minima, AMP optimally estimates the posterior marginals.
Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense). Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$. The fact that AMP follows state evolution is proven for such problems in [Javanmard and Montanari, 2013, Gerbelot and Berthier, 2023].
Appendix B Proofs of the main results
B.1 Linear Approximation
Our analysis relies on the following result:
Lemma B.1.
Let $F({{\bm{M}}})$ be as defined in Lemma 2.1
$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\mathcal{O}(\alpha\norm{\delta{{\bm{M}}}}_{F}^{2}),$ (30)
where $\norm{}_{F}$ denotes the Frobenius norm and $\mathcal{F}(\delta{{\bm{M}}})$ is a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$:
$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]\,,$ (31)
We proceed through an-entry-wise expansion of each term inside the expection in $F({{\bm{M}}})$ around ${{\bm{M}}}=0$. Recall that ${\bm{g}}_{\rm out}\in\mathcal{C}^{2}$ by Assumption 1. Since ${{\bm{M}}}_{F}\leq p$ , the first two derivatives of ${\bm{g}}_{\rm out}$ are uniformly bounded in ${\bm{g}}_{\rm out}$ for any fixed $Y,{\bm{\xi}}$. Therefore, applying the multivariate Taylor expansion to ${\bm{g}}_{\rm out}(Y^{t},\Delta_{1}{\bm{\xi}},I+\Delta_{2})$, with $\Delta_{1}=\sqrt{M}$ and $\Delta_{2}=,{\bm{I}}_{p}-{{\bm{M}}}$ yields:
${\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}$
$={\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}+\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\sqrt{M}{\bm{\xi}}{\bm{\xi}}^{\top}\sqrt{M}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}$
$+\langle\partial_{{\bm{V}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)M\rangle{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}+{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\langle M,\partial_{{\bm{V}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\rangle+\alpha C({\bm{\xi}},y)\mathcal{O}(\norm{M}^{2}_{F}),$
where $C({\bm{\xi}},Y)$ is an integrable function in ${\bm{\xi}},Y$. Since $T^{\star}$ is empty, ${\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)$ vanishes almost surely over $y$. Therefore, using dominated-convergence theorem, we obtain:
$F({{\bm{M}}})$ $=\alpha\mathbb{E}\left[{\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right){\bm{g}}_{\rm out}\!\left(Y^{t},\sqrt{{{\bm{M}}}^{t}}{\bm{\xi}},{\bm{I}}_{p}-{{\bm{M}}}^{t}\right)^{\top}\right]$
$=\alpha\mathbb{E}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)\sqrt{M}{\bm{\xi}}{\bm{\xi}}^{\top}\sqrt{M}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}\right]+\alpha\mathcal{O}(\norm{M}^{2}_{F})$
Since at ${{\bm{M}}}^{t}=0$, ${\bm{\xi}}$ and $Y$ are independent, the above simplifies to:
$F({{\bm{M}}})$ $=\alpha\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]+\alpha\mathcal{O}(\norm{M}^{2}_{F})$
B.2 Proof of Lemma 4.1
Suppose that $M\in\mathcal{S}^{+}$, we have, for any $\mathbf{v}\in\mathbb{R}^{p}$:
$\mathbf{v}^{\top}\mathcal{F}({{\bm{M}}})\mathbf{v}=\alpha\,\mathbb{E}_{y}\left[\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\mathbf{v}\right]\,,$ (32)
since ${{\bm{M}}}\in\mathcal{S}^{+}$, each term inside the expectation is non-negative. Therefore:
$\mathbf{v}^{\top}\mathcal{F}({{\bm{M}}})\mathbf{v}\geq 0$ (33)
Thus, $\mathcal{F}(M)$ is a cone-preserving linear map. The generalised Perron-Frobenius theorem for cone-preserving maps [Krein and Rutman, 1948] then implies that the operator $\mathcal{F}({{\bm{M}}})$ admits a unique eigenvector $M^{*}\in\mathcal{S}^{+}_{p}$ corresponding to the largest eigenvalue $\lambda_{\mathcal{F}}$ such that for any $M\in\mathcal{S}^{+}_{p}\perp M^{*}$:
$\mathcal{F}({{\bm{M}}})<\lambda_{\mathcal{F}}\norm{{{\bm{M}}}}_{F}.$ (34)
Furthermore, all other eigenvalues are strictly smaller than $\lambda_{\mathcal{F}}$. Subsequently, Lemma B.1 implies that $F({{\bm{M}}})$ is stable at ${{\bm{M}}}=0$ if and only if $\alpha\leq\frac{1}{\lambda_{\mathcal{F}}}=\alpha_{c}$.
B.3 Proof of Theorem 4.2
Lemma 2.1 allows us to map the behavior of the variable ${{\bm{M}}}^{t}$ in state-evolution to high-probability statements for the limiting overlaps Suppose that $A\in\mathcal{S}^{+}_{p}$ is a full rank matrix as in Theorem 4.2. Applying Lemmas B.1 and 3.2, we obtain that for ${{\bm{M}}}^{0}=\epsilon A$:
$\norm{{{\bm{M}}}^{t+1}}_{F}\leq\alpha\lambda_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F},$ (35)
for some constant $C$. Now, suppose that $\alpha<\frac{1}{\lambda_{F}}$ and $\epsilon<\frac{1}{C}(1-\alpha\lambda_{\mathcal{F}}-\kappa)$, for some $0<\kappa<1-\alpha\lambda_{\mathcal{F}}$ then inductively, we obtain:
$\norm{{{\bm{M}}}^{t+1}}_{F}<(1-\kappa)\norm{{{\bm{M}}}^{t}}_{F},$ (36)
for all $t>0$. Implying that $\norm{{{\bm{M}}}^{t}}_{F}\rightarrow 0$.
Now, suppose that $\alpha>\alpha_{c}$ or equivalently $\alpha\lambda_{F}>1$. Since $A$ is full-rank admits a decomposition:
$\operatorname{tr}(A,{{\bm{M}}}^{\star})=\theta$ (37)
for some $\theta>0$. From Lemma 4.1 and B.1, we obtain:
$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq\alpha\lambda_{\mathcal{F}}\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (38)
For $\norm{{{\bm{M}}}^{t}}\leq\frac{\theta}{C}(\alpha\lambda_{\mathcal{F}}-1-\kappa)$ for some $\kappa>0$, we inductively obtain:
$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq(1+\kappa)\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star}),$ (39)
implying that $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})$ grows as $\omega(e^{\kappa t})$ for sufficiently small $\epsilon$.
Finally, it remains to show that for large enough $\alpha$, ${{\bm{M}}}^{t}$ spans $E^{\star}$.
Notice that for any $\mathbf{v}\in E^{\star}$:
$\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v=\mathbb{E}_{y}\left[(\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\mathbf{v})^{2}\right]>0,$ (40)
while for any $A\in\mathcal{S}^{+}_{p}$:
$\mathbf{v}^{\top}\mathcal{F}(A)v\geq 0,$ (41)
since $\mathcal{F}(A)\in\mathcal{S}^{+}_{p}$. Define:
$\nu_{\mathcal{F}}=\inf_{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (42)
Since $\mathbf{v}$ lies in a compact set, $\nu_{\mathcal{F}}>0$. Let $A^{\star}$ denote the projection of $A$ along $E^{\star}$ and consider its eigendecomposition $A^{\star}$:
$A^{\star}=\sum_{i=1}^{p}\lambda_{i}\mathbf{v}_{i}\mathbf{v}^{\top}_{i}$ (43)
We
$\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (44)
Therefore, for $\alpha\geq\frac{1}{\nu_{\mathcal{F}}}$ and small enough $\epsilon$, ${{\bm{M}}}^{t}$ expands linearly along each $\mathbf{v}_{i}$.
B.4 Proof of Lemma 4.3
From the definition of $g_{out}$ in Equation 4 and the dominated-convergence theorem, it is straightforward to see that:
$\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}}=\mathbf{0},{\bm{I}})=\mathbb{E}[{\bm{Z}}{\bm{Z}}^{\top}-{\bm{I}}|Y].$ (45)
Therefore for any $\mathbf{v}\in\mathbb{R}^{p}$ with $\norm{\mathbf{v}}=1$:
$\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(Y,{\bm{\omega}},{\bm{I}})_{{\bm{\omega}}=\mathbf{0}}\mathbf{v}$ $=\mathbb{E}[(\mathbf{v}^{\top}{\bm{Z}})^{2}-1|Y]$
$=\lim_{d\rightarrow\infty}\mathbb{E}[\rm He_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star}\mathbf{x}\rangle)|y=Y],$
where the last expectation is w.r.t the joint measure $p(\mathbf{x},y)$ defined in Definition 1.
Now, for any measurable transformation $\mathcal{T}:\mathbb{R}\rightarrow\mathbb{R}$, we have, by the tower law of expectation:
$\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\rm He_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)|y\right]\right].$ (46)
For any $\mathbf{v}\in H^{\star}_{E}$, the integrand vanishes asymptotically almost surely over $y$. Therefore, the dominated-convergence theorem implies that for $\mathbf{v}\in H^{\star}_{E}$:
$\lim_{d\rightarrow\infty}\mathbb{E}\left[\mathcal{T}(y){\rm He}_{2}(\langle\mathbf{v}^{\top}{{\bm{{W}}}}^{\star},\mathbf{x}\rangle)\right]=0.$ (47)
B.5 Proof of Theorem 5.1
By assumption, $M_{U}$ is a fixed point of $F(M)$. Therefore, Equation 7 implies that ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}-\sqrt{M}_{U})$ almost surely lies in $U$ and thus $T^{\star}_{U}$ is empty. The proof of Theorem 5.1 then follows that of 4.2 by considering the following linearized operator along $U^{\perp}$
$\mathcal{F}_{{{\bm{M}}}_{U}}({{\bm{M}}}_{\perp})=\mathrel{\mathop{:}}=\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U}){{\bm{M}}}_{\perp}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,\sqrt{{{\bm{M}}}_{U}}{\bm{\xi}},{\bm{I}}-\sqrt{M}_{U})^{\top}\right]\,,$ (48)
where ${\bm{\omega}}=\sqrt{M}_{U}{\bm{\xi}}$ with ${\bm{\xi}}\sim\mathcal{N}(0,{\bm{I}})$. Similar to Lemma B.1, the linearization follows by noting that $y$ is independent of $\sqrt{{{\bm{M}}}_{\perp}}{\bm{\xi}}$ and $\sqrt{{{\bm{M}}}_{U}+{{\bm{M}}}_{\perp}}=\sqrt{{{\bm{M}}}_{U}}+\sqrt{{{\bm{M}}}_{\perp}}$ since ${{\bm{M}}}_{\perp}\perp{{\bm{M}}}_{U}$.
Appendix C A list of examples of the ${\bm{g}}_{\rm out}$ function
In this section, we list several problems, their corresponding ${\bm{g}}_{\rm out}$, and analysis of their weak recoverability.
C.1 $g(z_{1},z_{2},....,z_{p})=\prod_{j=1}^{p}z_{j}$
For all $p$, ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=0$. Otherwise we have two cases We have two cases. If $p=2$:
$\mathcal{Z}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=\frac{K_{0}(|y|)}{\pi}$ (49)
and
$\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=\begin{bmatrix}|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}-1&y\\ y&|y|\frac{K_{1}(|y|)}{K_{0}(|y|)}-1\\ \end{bmatrix}$ (50)
where $K_{n}(y)$ are the modified Bessel functions of the second kind. If $p>3$ then
$\mathcal{Z}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=\frac{1}{(2\pi)^{p/2}}G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right)$ (51)
The matrix $\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})$ is diagonal, with identical diagonal elements equal to $f(y)$:
$f(y)=2G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0\end{array}\right)/G^{p,0}_{0,p}\left(\frac{y^{2}}{2^{p}}\,\bigg{|}\,\begin{array}[]{c}0\\ 0,0,\ldots,0,1\end{array}\right)-1$ (52)
The alpha critical $\alpha_{c}$ will thus be
$\alpha_{c}=\left[\int_{-\infty}^{\infty}\,{\rm d}yf(y)^{2}\mathcal{Z}_{\rm out}(y)\right]^{-1}$ (53)
C.2 $\frac{1}{p}\sum_{i=1}^{p}z_{i}^{2}$
The function is even, so ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})=0$. The trick for the computation is to move to generalised spherical coordinates by choosing $r^{2}=z_{2}^{2}+...z_{p}^{2}$. Recall that the area of the unit sphere in $p-1$ dimensions is
$\frac{2\pi^{\frac{p-1}{2}}}{\Gamma\left(\frac{p-1}{2}\right)}$ (54)
The matrix $\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}_{p})$ is diagonal because of parity. All the elements on the diagonal are
$(y-1)^{2}\frac{2^{-\frac{p}{2}}e^{-\frac{py}{2}}(py)^{\frac{p}{2}}}{\Gamma\left(\frac{p}{2}\right)}$ (55)
The critical alpha will be
$\alpha_{c}=\left[\int\,{\rm d}y(y-1)^{2}\frac{2^{-\frac{p}{2}}e^{-\frac{py}{2}}(py)^{\frac{p}{2}}}{\Gamma\left(\frac{p}{2}\right)}\right]^{-1}=\frac{p}{2}$ (56)
Of course this procedure fail at the first step if $p=1$. On the other hand, one can check that the final result is still correct for all positive integer $p$.
C.3 $g(z_{1},\ldots,z_{p})=\operatorname{sign}(\prod_{j=1}^{p}z_{j})$
We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
$d_{ij}({\bm{V}})$ $:=\det\left(\begin{array}[]{cc}v_{i,j}&v_{i,p}\\ v_{j,p}&v_{p,p}\end{array}\right),\quad i,j\in{1,\ldots,p},$ (59)
the function
$E_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}})=1-s\operatorname{erf}\left(\frac{v_{1,p}(\sum_{j=1}^{p-1}v_{j,p}(z_{j}-\omega_{j})-v_{pp}\omega_{p}}{\sqrt{2v_{pp}}}\right),$
the Gaussian probability density
$\rho({\bm{z}}\in\mathbb{R}^{p},{\bm{V}}):=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}\sqrt{\det({\bm{V}})}}e^{-\frac{1}{2}{\bm{z}}^{\top}{\bm{V}}^{-1}{\bm{z}}},$ (60)
and finally
$I_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}}):=E_{s}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\rho\left(\left(\begin{array}[]{ccc}z_{1}-\omega_{2}\\ \ldots\\ z_{p-1}-\omega_{p-1}\end{array}\right)^{\top},v_{pp}\bm{d}({\bm{V}})^{-1}\right).$
Then, defining $\mathbb{R}_{+1}=[0,+\infty)$ and $\mathbb{R}_{-1}=(-\infty,0]$, we have that
$Z_{\rm out}(y,{\bm{\omega}}{\bm{V}})=\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{\delta_{y,(\prod_{j}s_{j})}}{2}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}}).$ (61)
The components of ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})$ are given by
$({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}}))_{k}$ $=\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{s_{k}\delta_{y,(\prod_{j}s_{j})}}{2}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
Stability of the fixed point
We can now focus on studying the stability
$Z_{\rm out}(y,\bm{0},{\bm{I}})=\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=\frac{1}{2}$ (62)
In order to compute $\nabla_{\omega}{\bm{g}}_{\rm out}(y,\bm{0},{\bm{I}})$, we note at first that
$\frac{1}{(2\pi)^{p/2}}\int{\rm d}{\bm{z}}(z_{k}^{2}-1)e^{-\nicefrac{{{\bm{z}}^{\top}{\bm{z}}}}{{2}}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}=0,\quad\forall k\in\{1,\ldots,p\}.$ (63)
For $p=2$, consider the integral
$\frac{1}{2\pi}\int{\rm d}z_{1}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}\delta_{y,\operatorname{sign}(z_{1}\ldots z_{p})}$ (64)
$=$ $\frac{2}{2\pi}\left(\delta_{y,1}\int_{0}^{\infty}{\rm d}z_{1}\int_{0}^{\infty}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}+\delta_{y,-1}\int_{0}^{\infty}{\rm d}z_{1}\int_{-\infty}^{0}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}\right)$ (65)
$=\frac{{\rm sign}{(y)}}{\pi}$ (66)
Using (17) we obtain that $\alpha_{c}$ is the inverse of the largest eigenvalues of
$\frac{4}{\pi^{2}}\begin{bmatrix}0&1\\ 1&0\\ \end{bmatrix}$ (67)
So $\alpha_{c}=\pi^{2}/4$. For $p\geq 3$, defining (±)$\mathbb{R}^{p}_{(\pm)}=\{{\bm{z}}\in\mathbb{R}^{p}|\operatorname{sign}(z_{1}\ldots z_{p})=\pm 1\}$, we need to compute integrals of the type
$\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int_{\mathbb{R}^{p}_{(+)}}{\rm d}z_{1}\ldots{\rm d}z_{p}z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}}\ldots e^{-\frac{z_{p}^{2}}{2}}$
$=$ $\frac{2}{2^{p-2}(2\pi)^{\nicefrac{{p}}{{2}}}}\left(\sum_{j=0}^{\lfloor\frac{p-2}{2}\rfloor}\left(\begin{array}[]{cc}p-2\\ 2j\end{array}\right)\int_{\mathbb{R}^{2}_{(+)}}{\rm d}z_{1}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}+\sum_{j=1}^{\lfloor\frac{p-1}{2}\rfloor}\left(\begin{array}[]{cc}p-2\\ 2j-1\end{array}\right)\int_{\mathbb{R}^{2}_{(-)}}{\rm d}z_{1}{\rm d}z_{2}\,z_{1}z_{2}e^{-\frac{z_{1}^{2}}{2}-\frac{z_{2}^{2}}{2}}\right)$
$=$ $\frac{2^{3-p}}{(2\pi)^{(p+2)/2}}\sum_{j=0}^{p-2}\left(\begin{array}[]{cc}p-2\\ j\end{array}\right)(-1)^{j}=0$
In the same way it is possible to prove that the integral over $\mathbb{R}_{(-)}^{p}$ is also vanishing. This implies that $\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}=0$, so using (17), we obtain that $\alpha_{c}=+\infty$. This model cannot be learned with $n=\mathcal{O}(d)$ samples for $p\geq 3$.
C.4 $g(z_{1},\ldots,z_{p})=z_{1}^{2}+\operatorname{sign}\left(\prod_{j=1}^{p}z_{j}\right)$
We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
$d_{ij}({\bm{V}})$ $:=\det\left(\begin{array}[]{cc}v_{i,j}&v_{i,p}\\ v_{j,p}&v_{p,p}\end{array}\right),\quad i,j\in{1,\ldots,p},$ (70)
the function
$E_{s,s_{1},s_{p}}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}})=1-s_{p}\operatorname{erf}\left(\frac{v_{1,p}(s_{1}\sqrt{y-s}-\omega_{1})+\sum_{j=2}^{p-1}v_{j,p}(z_{j}-\omega_{j})-v_{pp}\omega_{p}}{\sqrt{2v_{pp}}}\right),$
the Gaussian probability density
$\rho({\bm{z}}\in\mathbb{R}^{p},{\bm{V}}):=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}\sqrt{\det({\bm{V}})}}e^{-\frac{1}{2}{\bm{z}}^{\top}{\bm{V}}^{-1}{\bm{z}}},$ (71)
and finally
$I_{s,s_{1},s_{p-1}}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}}):=\frac{E_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})}{4\sqrt{y-s}}\rho\left(\left(\begin{array}[]{cccc}s_{1}\sqrt{y-s}-\omega_{1}\\ z_{2}-\omega_{2}\ldots\\ z_{p-1}-\omega_{p-1}\end{array}\right),v_{pp}\bm{d}({\bm{V}})^{-1}\right).$
Then, defining $\mathbb{R}_{+1}=[0,+\infty)$ and $\mathbb{R}_{-1}=(-\infty,0]$, we have that
$Z_{\rm out}(y,{\bm{\omega}}{\bm{V}})=\sum_{s,s_{1},\ldots,s_{p}=\pm 1}\delta_{s,(\prod_{j}s_{j})}\bm{1}_{y>s}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}}),$ (72)
where $\bm{1}_{x>0}$ is the Heaviside step function. The components of ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})$ are given by
$({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}}))_{1}$ $=\sum_{s,s_{1},\ldots,s_{p}=\pm 1}\frac{\delta_{s,(\prod_{j}s_{j})}\bm{1}_{y>s}}{Z_{\rm out}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}\frac{\partial}{\partial\omega_{1}}I_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})$
$({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}}))_{k\neq 1}$ $=\sum_{s,s_{1},\ldots,s_{p}=\pm 1}\frac{s_{k}\delta_{s,(\prod_{j}s_{j})}\bm{1}_{y>s}}{Z_{\rm out}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s,s_{1},s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
Stability of the fixed point
The result is dependent on a number of integrals which we can compute analytically. First we have
$Z_{\rm out}(y,\bm{0},{\bm{I}})$ $=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (73)
$=\frac{1}{2(2\pi)^{\nicefrac{{p}}{{2}}}}\sum_{s_{1},\ldots,s_{p}=\pm 1}\frac{\bm{1}_{y>s}}{\sqrt{y-\prod_{j}s_{j}}}\int_{\mathbb{R}_{s_{2}}}{\rm d}z_{2}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}e^{-\frac{y-s_{1}\ldots s_{p}}{2}-\frac{z_{2}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}$ (74)
$=\frac{1}{2\sqrt{2\pi}}\sum_{s=\pm 1}\bm{1}_{y>s}\frac{e^{-\frac{y-s}{2}}}{\sqrt{y-s}}$ (75)
$=\frac{e^{-\nicefrac{{y}}{{2}}}}{2\sqrt{2\pi e}}\left(\frac{1}{\sqrt{y+1}}+\mathbf{1}_{y>1}\frac{e}{\sqrt{y-1}}\right)$ (76)
From this it is straightforward to see that ${\bm{g}}_{\rm out}^{0}(y)=0$. In order to compute $\nabla_{\omega}{\bm{g}}_{\rm out}(y,0,1)$ we need to consider the additional integrals
1 1)\;1 ) $\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}(z_{1}^{2}-1)e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))$ (77)
$=\frac{1}{2\sqrt{2\pi}}\sum_{s=\pm 1}\bm{1}_{y>s}\frac{e^{-\frac{y-s}{2}}(y-s-1)}{\sqrt{y-s}}$ (78)
$=\frac{e^{-y}}{2\sqrt{2\pi e}}\left(\frac{y}{\sqrt{y+1}}+\mathbf{1}_{y>1}\frac{e(y-2)}{\sqrt{y-1}}\right)$ (79)
2 2)\;2 ) $\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}(z_{k\neq 1}^{2}-1)e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))=0$ (80)
3 3)\;3 ) $\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}}\int{\rm d}z_{1}\ldots{\rm d}z_{p}z_{j}z_{k}e^{-\frac{z_{1}^{2}}{2}-\ldots-\frac{z_{p}^{2}}{2}}\delta(y-z_{1}^{2}-{\rm sign}(z_{1}\ldots z_{p}))$ (81)
$\propto\sum_{s,s_{j},s_{k}=\pm 1}\bm{1}_{y>s}s_{j}s_{k}e^{-\frac{y-s}{2}}=0$ (82)
This shows that
$\partial_{\bm{\omega}}{\bm{g}}_{\rm out}(y)=\begin{bmatrix}C(y)&0&0\\ 0&0&0\\ 0&0&0\\ \end{bmatrix}$ (84)
Where $C(y)$ is
$C(y)=\begin{cases}y&-1<y<1\\ y-2e\frac{\sqrt{y+1}}{e\sqrt{y+1}+\sqrt{y-1}}-1&y>1\\ \end{cases}$ (85)
The critical alpha is simply found by one last integral
$\alpha_{c}=\left[\int\,{\rm d}yC(y)^{2}Z_{\rm out}(y)\right]^{-1}\approx 0.575166$ (86)
C.5 $g(z_{1},\ldots,z_{p})=\sum_{j=1}^{p}\operatorname{sign}(z_{p})$
We define ${\bm{\omega}}=(\omega_{1},\ldots,\omega_{p})^{\top}$ and $v_{ij}\coloneqq({\bm{V}}^{-1})_{i,j}$. We also introduce the matrix $\bm{d}({\bm{V}})\in\mathbb{R}^{(p-1)\times(p-1)}$ such that
$d_{ij}({\bm{V}})$ $:=\det\left(\begin{array}[]{cc}v_{i,j}&v_{i,p}\\ v_{j,p}&v_{p,p}\end{array}\right),\quad i,j\in{1,\ldots,p},$ (89)
the function
$E_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}})=1-s\operatorname{erf}\left(\frac{\sum_{j=1}^{p-1}v_{j,p}(z_{j}-\omega_{j})-v_{pp}\omega_{p}}{\sqrt{2v_{pp}}}\right),$ (90)
the Gaussian probability density
$\rho({\bm{z}}\in\mathbb{R}^{p},{\bm{V}}):=\frac{1}{(2\pi)^{\nicefrac{{p}}{{2}}}\sqrt{\det({\bm{V}})}}e^{-\frac{1}{2}{\bm{z}}^{\top}{\bm{V}}^{-1}{\bm{z}}},$ (91)
and finally
$I_{s}({\bm{z}}\in\mathbb{R}^{p-1},y,{\bm{\omega}},{\bm{V}}):=E_{s}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\rho\left(\left(\begin{array}[]{ccc}z_{1}-\omega_{2}\\ \ldots\\ z_{p-1}-\omega_{p-1}\end{array}\right)^{\top},v_{pp}\bm{d}({\bm{V}})^{-1}\right).$
Then, defining $\mathbb{R}_{+1}=[0,+\infty)$, $\mathbb{R}_{-1}=(-\infty,0]$ and $S_{a}=\{\bm{s}=(s_{1},...,s_{p})|\exists\bm{j}=\{j_{1},...,j_{a}\}\subseteq\{1,...,n\}\text{ s.t }s_{j}=1\text{ for }j\in\bm{j}\text{ and }s_{j}=-1\text{ otherwise}\}$, we have that
$Z_{\rm out}=\sum_{a=1}^{p}\frac{\delta_{y,2a-p}}{2}\sum_{\bm{s}\in S_{a}}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}}),$ (92)
and the $k^{\text{th}}$ component of ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})$ is
$\left({\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{V}})\right)_{k}=\frac{1}{Z_{\rm out}}\sum_{a=1}^{p}\frac{\delta_{y,2a-p}}{2}\sum_{\bm{s}\in S_{a}}s_{k}\int_{\mathbb{R}_{s_{1}}}{\rm d}z_{1}\ldots\int_{\mathbb{R}_{s_{p-1}}}{\rm d}z_{p-1}I_{s_{p}}({\bm{z}},y,{\bm{\omega}},{\bm{V}})\delta(z_{k}-0)$
Appendix D Further numerical observations and details
Here we give more details about the numerical implementation of State Evolution and AMP. Both approaches require to compute ${\bm{g}}_{\rm out}$. All the examples we implemented are detailed in Appendix C.
The integrals are doing with the quadrature package in Scipy. In order to avoid instabilities we regularize the interval of integration by replacing $\infty$ with $\Lambda$. We typically choose $\Lambda\approx 10$. Similarly, we add $\epsilon\approx 10^{-4}$ to the diagonal of $V$. In State Evolution we need to integrate over functions of ${\bm{g}}_{\rm out}$. We do these integrals using a simple Monte Carlo approach. For Figures 1 and 2 we used a total $72000$ samples, for Figure 4 $7200$. Computing such integrals is the numerical bottleneck. In order to make this part faster we parallelised the MCMC: for each iteration of State Evolution we make every worker in our pool estimate the integral, and then average the estimation and then average among workers. In the cases in which ${{\bm{M}}}=0$ is a fixed point, we initialize ${{\bm{M}}}$ with the empirical overlap of AMP at the beginning of the iteration.
In both the AMP and State Evolution implementation we used some damping: the overlap ${{\bm{M}}}$ or the $\hat{{{\bm{{W}}}}}$ at the new iterations are averaged with the current value, with a weight $\delta$ for the new one, where typically $0.90.6<\delta<$ 0.9. We display the evolution of the overlaps in a typical run of AMP for the model $z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$ in Figure 3. We already displayed the values of the overlap and generalisation error at convergence in Figure 2. We can see how AMP has a saddle-to-saddle dynamic, where the algorithm alternates plateaus for $\mathcal{O}(\log{d})$ ) iterations to fast drops in generalisation error, which are associated with new directions being learned.
As stated in the main text, models non-trivial subspaces are associated with symmetries. This also implies that the associated overlaps are have the same invariances. In order to make the plots readable we remove all such symmetries by hand. For this reason we take the absolute value of the overlap if the model is even, and we impose a specific inequality in the overlap if there is invariance under permutation. As a general idea we want to always have the ”best” possible overlap. Meaning we want ${{\bm{M}}}$ to be as diagonal as possible. We list what this mean for the examples in the figures:
  • ${\rm sign}(z_{1}z_{2})$: Because of invariance under exchange of $z_{1}$ and $z_{2}$ we always have ${{\bm{M}}}_{11}={{\bm{M}}}_{22}$. Here AMP can reach $2$ equivalent configurations: either the diagonal or the anti-diagonal is zero. We choose the configuration where the anti-diagonal is zero.
  • $z_{1}^{2}+{\rm sign}(z_{1}z_{2}z_{3})$: This model will first learn just $z_{1}$. We fix ${{\bm{M}}}_{11}>0$. For the rest of the components we are reduced to the case above.
  • ${\rm sign}(z_{1})+{\rm sign}(z_{2})+{\rm sign}(z_{3})$: Because of the invariance under permutation each row of ${{\bm{M}}}$ will have either the same element in all the entries (in which case we don’t need to do anything) or two are the same and one is bigger. We permute the matrix such that the largest entry is on the diagonal.
The code to run AMP and State Evolution on the examples in the is available on GitHub